# Right Time to Learn: Promoting Generalization via Bio-inspired Spacing Effect in Knowledge Distillation

### v1

@@ Abstract
Knowledge distillation (KD) is a powerful strategy for training deep neural networks (DNNs). Although it was originally proposed to train a more compact “student” model from a large “teacher” model, many recent efforts have focused on adapting it to promote generalization of the model itself, such as online KD and self KD. Here, we propose an accessible and compatible strategy named Spaced KD to improve the effectiveness of both online KD and self KD, in which the student model distills knowledge from a teacher model trained with a space interval ahead. This strategy is inspired by a prominent theory named spacing effect in biological learning and memory, positing that appropriate intervals between learning trials can significantly enhance learning performance. With both theoretical and empirical analyses, we demonstrate that the benefits of the proposed Spaced KD stem from convergence to a flatter loss landscape during stochastic gradient descent (SGD). We perform extensive experiments to validate the effectiveness of Spaced KD in improving the learning performance of DNNs (e.g., the performance gain is up to 2.31% and 3.34% on Tiny-ImageNet over online KD and self KD, respectively).
@@ 3.1 Problem Setup
The design of an online teacher is quite demand-oriented, it could be simply a copy of the student network (Li et al., 2022b; Wu & Gong, 2021). But to maintain a valid knowledge gap between student and teacher, they are often initialized using different random seeds in practice. Besides, the training process of teacher network could also be intervened by auxiliary loss from students through reverse distillation (Li & Jin, 2022; Qian et al., 2022; Shi et al., 2021).
@@ 5.1 Experimental Setups
For ResNet-based architectures, we use an SGD optimizer (Sutskever et al., 2013) with 0.9 momentum, 128 batch size, 80 epochs, and a constant learning rate of 0.01. For KD-related hyperparameters (Zhang et al., 2019), we use a distillation temperature of 3.0, a feature loss coefficient of 0.03, and a KL-Divergence loss weight of 0.3. For transformer-based architectures, we use an AdamW optimizer (Loshchilov & Hutter, 2017a) of batch size 128 and epoch number 300 (warm-up for 20 epochs). Besides, a cosine learning rate decay policy (Loshchilov & Hutter, 2017b) is utilized with initial learning rate $45e-$ 4 and final $65e-$ 6, following the training pipeline of previous works (Liu et al., 2021; Li et al., 2022a; Sun et al., 2024).
@@ 5.2 Effectiveness and Generality of Spacing Effect
We combine our Spaced KD with other more advanced KD methods, including (1) traditional KD such as BAN (Furlanello et al., 2018) and TAKD (Mirzadeh et al., 2020a), (2) online KD such as DML (Zhang et al., 2018) and SHAKE (Li & Jin, 2022), and (3) self KD such as DLB (Shen et al., 2022) and PSKD (Kim et al., 2021) (see Appendix A.3 for details). As shown in Tab. 4, Spaced KD brings significant improvements to a wide range of KD methods. The above results suggest that the benefits of Spaced KD arise from the fundamental properties of parameter optimization in deep learning, consistent with our theoretical analysis in Sec. 4.4.
@@
Method w/o KD w/ KD w/ Ours
BAN (Furlanello et al., 2018) 56.75 60.56 61.83
TAKD (Mirzadeh et al., 2020a) 61.37 61.82 63.48
DML (Zhang et al., 2018) 68.92 70.31 71.80
SHAKE (Li & Jin, 2022) 69.02 72.02 72.64
DLB (Shen et al., 2022) 68.80 68.87 69.31
PSKD (Kim et al., 2021) 74.92 75.20 75.38
@@ 5.3 Extended Analysis of Spacing Effect
As described in previous works, the learning rate and batch size influence the endpoint curvature and the whole trajectory (Frankle et al., ; Lewkowycz et al., 2020; Xie et al., 2020). The learning rate corresponds to the parameters’ updating step length, and batch size would affect the total number of updating iterations which directly relates to the choice of space interval $s$. Therefore, we further validate the impact of learning rate and batch size. As shown in Fig. 5 of Appendix A.7, we summarize the results: (i) Spaced KD proves effective w.r.t naive online KD across different learning rates; (ii) Spaced KD exhibits its advantages when training with a relatively large batch size (greater than 64). These observations also align with previous research (Jastrzebski et al., 2019; Wu et al., ) regarding a small batch size limiting the maximum spectral norm along the convergence path found by SGD from the beginning of training.
@@
Attack ResNet-18 ResNet-50 ResNet-101

### v2

@@ Abstract
Knowledge distillation (KD) is a powerful strategy for training deep neural networks (DNNs). Although it was originally proposed to train a more compact “student” model from a large “teacher” model, many recent efforts have focused on adapting it to promote generalization of the model itself, such as online KD and self KD. Here, we propose an accessible and compatible strategy named Spaced KD to improve the effectiveness of both online KD and self KD, in which the student model distills knowledge from a teacher model trained with a space interval ahead. This strategy is inspired by a prominent theory named spacing effect in biological learning and memory, positing that appropriate intervals between learning trials can significantly enhance learning performance. With both theoretical and empirical analyses, we demonstrate that the benefits of the proposed Spaced KD stem from convergence to a flatter loss landscape during stochastic gradient descent (SGD). We perform extensive experiments to validate the effectiveness of Spaced KD in improving the learning performance of DNNs (e.g., the performance gain is up to 2.31% and 3.34% on Tiny-ImageNet over online KD and self KD, respectively). Our codes have been released on github https://github.com/SunGL001/Spaced-KD.
@@ 3.1 Problem Setup
The design of an online teacher is quite demand-oriented, it could be simply a copy of the student network (Li et al., 2022c; Wu & Gong, 2021). But to maintain a valid knowledge gap between student and teacher, they are often initialized using different random seeds in practice. Besides, the training process of teacher network could also be intervened by auxiliary loss from students through reverse distillation (Li & Jin, 2022; Qian et al., 2022; Shi et al., 2021).
@@ 5.1 Experimental Setups
For ResNet-based architectures, we use an SGD optimizer (Sutskever et al., 2013) with 0.9 momentum, 128 batch size, 80 epochs, and a constant learning rate of 0.01. For KD-related hyperparameters (Zhang et al., 2019), we use a distillation temperature of 3.0, a feature loss coefficient of 0.03, and a KL-Divergence loss weight of 0.3. For transformer-based architectures, we use an AdamW optimizer (Loshchilov & Hutter, 2017a) of batch size 128 and epoch number 300 (warm-up for 20 epochs). Besides, a cosine learning rate decay policy (Loshchilov & Hutter, 2017b) is utilized with initial learning rate $45e-$ 4 and final $65e-$ 6, following the training pipeline of previous works (Liu et al., 2021; Li et al., 2022b; Sun et al., 2024).
@@ 5.2 Effectiveness and Generality of Spacing Effect
We combine our Spaced KD with other more advanced KD methods, including (1) TSB (Li et al., 2022a) constructs superior ”teachers” with temporal accumulator and spatial integrator; (2) CTKD (Li et al., 2023) controls the task difficulty level during the student’s learning career through a dynamic and learnable temperature; (3) LSKD (Sun et al., 2024) employs a plug-and-play Z-score pre-process of logit standardization before applying softmax and KL divergence. As shown in Tab. 4, Spaced KD brings significant improvements to a wide range of KD methods. The above results suggest that the benefits of Spaced KD arise from the fundamental properties of parameter optimization in deep learning, consistent with our theoretical analysis in Sec. 4.4.
@@
Dataset/Model Method TSB CTKD LSKD
CIFAR-100/ResNet-18 w/o KD 67.65 67.86 67.94
w/ KD 71.70 69.41 70.74
w/ Ours 72.82 71.12 71.76
CIFAR-100/DeiT-Tiny w/o KD 51.90 53.31 52.31
w/ KD 52.63 54.20 52.93
w/ Ours 55.47 54.72 53.83
Tiny-ImageNet/ResNet-18 w/o KD 55.21 53.03 54.05
w/ KD 59.92 58.78 59.30
w/ Ours 61.65 60.32 60.28
Tiny-ImageNet/DeiT-Tiny w/o KD 40.29 40.82 39.65
w/ KD 40.13 41.22 41.14
w/ Ours 43.36 41.60 41.48
@@ 5.3 Extended Analysis of Spacing Effect
As described in previous works, the learning rate and batch size influence the endpoint curvature and the whole trajectory (Frankle et al., ; Lewkowycz et al., 2020; Xie et al., 2020). The learning rate corresponds to the parameters’ updating step length, and batch size would affect the total number of updating iterations which directly relates to the choice of space interval $s$. Therefore, we further validate the impact of learning rate and batch size. As shown in Fig. 5 of Appendix A.7, we summarize the results: (i) Spaced KD proves effective w.r.t naive online KD and self KD across different learning rates; (ii) Spaced KD exhibits its advantages when training with a relatively large batch size (greater than 64). These observations also align with previous research (Jastrzebski et al., 2019; Wu et al., ) regarding a small batch size limiting the maximum spectral norm along the convergence path found by SGD from the beginning of training.
@@
Method Attack ResNet-18 ResNet-50 ResNet-101
@@
Self KD impulse_noise 50.65 60.57 9.92 62.18 71.57 9.39 59.33 68.78 9.45
zoom_blur 64.44 68.60 4.16 68.03 72.13 4.10 66.09 71.16 5.07
snow 61.30 66.14 4.84 64.72 69.65 4.93 64.03 68.76 4.73
frost 63.96 67.80 3.84 66.36 71.81 5.45 66.73 70.21 3.48
jpeg_compression 30.99 34.67 3.68 34.44 35.34 0.90 33.64 34.76 1.12
brightness 73.18 75.92 2.74 74.91 79.19 4.28 75.10 78.91 3.81
@@ 6 Conclusion
Acknowledgment.
This work was supported by the STI2030-Major Projects (2022ZD0204900 to Y.Z.), the NSFC Projects (Nos. 62406160 to L.W., 32021002 to Y.Z.), and National Science and Technology Major Project (2022ZD01163013 to B.L.). L.W. is also supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZB20230350 and the Shuimu Tsinghua Scholar.